OpenStudy (anonymous):

A bag contains a total of 8 batteries, of which three are defective. Selecting three at random, without replacement, determine the probability that none of the batteries you select are good.

OpenStudy (anonymous):

if ur drawing the batteries at the same time , then the probability is 3/8 but wen ur drawing batteries one by one then, the probability is (3/8)*(2/7)8(1/6)

OpenStudy (anonymous):

gramiof2:got it?

OpenStudy (anonymous):

i think the 8 is a typo yes?

OpenStudy (anonymous):

\[\frac{3}{8}\times \frac{2}{7}\times \frac{1}{6}\]

OpenStudy (anonymous):

I think it is 1/56

OpenStudy (anonymous):

yea the 8 there was *---missed the shift key :)